The three capacitors C= 6.0 uF (micro F) in series are connected across a 30-V battery.

A). Find the charge on each capacitor.
B). Find the potential difference across each.
Answer #1

1/(1/6+1/6+1/6)≈2 μF=Ceq==>Q=C*V=30 V*2 μF=60 μC

A) 60 μC on each capacitor

B) 60 μC/6 μF=10 V across each capacitor